OpenStudy (anonymous):

find the equation of plane passing through point(0,7,-7) and containing line x+1/-3=y-3/2=z+2/1

OpenStudy (shubhamsrg):

Well I am not very good at 3D but I can probably tell you a method, which is lengthy, but you might get to the answer.

OpenStudy (shubhamsrg):

Let eqn of plane be Ax +By + Cz = D it passes through (0,7,-7) this gives you 1st eqn

OpenStudy (shubhamsrg):

Now it also contains the given line, now, as according to your liking, take any 3 random points lying on the line and substitute it in the eqn of the plane Thus, you'll get 4 eqns with 4 variables, solve for A,B,C and D This should help maybe ? I am not too sure.

OpenStudy (anonymous):

Let the direction ratios of the plane be a , b , c. direction ratio of given line -3,2,1 Now since the plane contain the line -3a+2b+c=0 __(1) equation of plane a(x-0) +b(y -7) +c(z-(-7)) =0 , now the point -1,3,-2 satisfy the equ. of plane... a(-1-0) +b(3-7) +c(-2+7)=0 -a-4b+5c=0 , a+4b-5c=0 _____(2) solving (1) & (2) we get a=14 , b=14 & c=14 hence equation of plane 14(x-0) + 14(y-7) +14(z+7)=0 x+y+z=0

OpenStudy (anonymous):

plse see , if i had done correct